OpenStudy (anonymous):

derivative of〖sec(〗⁡〖1/2∅)〗 tan⁡〖1/2∅〗

OpenStudy (ikram002p):

:O would u like to make effort and write it again :P with latex or draw tool ? \(\downarrow \)

OpenStudy (gudden):

@turnback .... kindly turn back to the question :)

OpenStudy (anonymous):

\[lety=\sec(1/2)\theta\tan(1/2)\theta\] \[\frac{ d }{ dx }\left( uv \right)=uv \prime+u \prime v\]

OpenStudy (anonymous):

here find \[d/ d\theta\]